OpenStudy (anonymous):

Suppose $800 is invested into an account that pays 3.5% interest, compounded annually. What is the balance in the account after 5 years? A. $560.00 B. $950.15 C. $1142.86 D. $2800.00

OpenStudy (anonymous):

$800 3.5 interest 5 years Total = Principal * (1+r)^years Total = 800 * (1.035)^5 Total = 800 * 1.1876863056 Total = ?????

OpenStudy (anonymous):

950.15 dollars

OpenStudy (anonymous):

yes!